# Consumption Smoothing in Metropolis: Evidence from Working-class Households in Prewar Tokyo

### v12

@@ B.2 Original Micro-spreadsheets from the Household Budget Survey in Tsukishima Survey
The official report does not include household-level information, and only a small quantity of the original forms used to prepare the report have survived. Fortunately, however, some of the original micro-spreadsheets are kept in the archives of the OISR because Takano was its first director. This study uses the original micro-spreadsheets from the household budget survey (No. 9 in Table B.1), which was used by Gonda to write the second section of the official report. While they have yet to be publicly available, I received permission from the OISR to use the original forms stored in the archives: the OISR, Archives of the Tsukishima Survey (THBS, unreleased). [^93] I would like to express my sincere thanks to Professor Kazue Enoki, Deputy Director of the OISR, for allowing me access to the documents. I wish to thank Ms. Mika Nakamura, an archivist at the institute, for her careful assistance during the research.
@@
Notes: This photograph depicts a sample page from the household expenses account of the THBS. The right page documents the details of costs. The left page includes the details of revenue and receiving and making payments on the goods (actual things, including gifts).
Source: The THBS, account #6, 15th February, 1919. The tone was adjusted by the author using Adobe Photoshop 24.7.0.
The budget book used for the household budget survey is named the Kinsendeiri hikaechō (money receipts and disbursements account book). The design of this budget book was based on the form used in Takano’s first household survey (Online Appendix B.1). [^94] Department of Health, Ministry of the Interior (1923a, p. 152). The budget book was to be filled out daily and included four sections: income, expenses, receipts from credit purchases, and receipts and disbursements of actual goods. Most of the entries were in the income and expenses sections, indicating that their daily living was conducted through wage receipt and consumption in cash. Figure B.3 shows an example of the budget book.
As described, the method used to select households for the survey was not recorded in detail (Online Appendix B.1). According to the report, they initially recruited just over $90$ households through the factories, police station, and elementary schools in Tsukishima. Of these, $50$ households completed the book for at least one month, but $10$ were deemed incomplete and not included in the report. In addition to the $40$ budget books used to prepare the report, therefore, the OISR has another set of budget books marked as “discarded.” According to my survey, about six incomplete books of the worker households have been determined to be in this “discarded” category. Therefore, based on the condition of the remaining books, it is plausible to believe that approximately $50$ households participated in the survey. Of the $40$ households included in the report tally, two households maintained records for more than one year, $11$ households for 6–12 months, and the remaining $27$ households registered for less than six months (Department of Health, Ministry of the Interior 1923a, pp. 152–153). Initially, the survey period was intended to be six months, but the average bookkeeping period for the $40$ households was four months.
@@ B.3 Trimming
(c) Adj. monthly panel dataset
@@
Note: Figure B.4aB.4b, and B.4c show the observations in the daily, semi-monthly, and adjusted monthly panel datasets, respectively. For example, a marker in Figure B.4a indicates an observation in a specific year-month-date cell. Strictly, semi-monthly, and monthly datasets include both panel and cross-sectional units because aggregation shall trim fractions in the edges of the survey period. Figure B.4b shows 35-panel units and five cross-sectional units (#29; #31-33; #35). Figure B.4c shows 28-panel units and six cross-sectional units (#20; #25; #26; #28; #34; #40). Five cross-sectional units in the semi-monthly panel do not appear in this figure (#29; #31-33; #35). Note that unit #30 having two semi-monthly cells is not included because those cells are separated into different adjusted months (February and March, respectively).
Source: Created by the author using the THBS sample.
To check whether the statistics from THBS budget books contained any missing information, I digitized all $40$ household budget books. Throughout this digitization process, I found that there are a few misreporting issues in the budget books.
First, there is a household that does not include income information (household #37). Although the official report by Gonda reported the average monthly income for this household (Department of Health, Ministry of the Interior 1923a, p. 183), no documentation enabled me to determine how this average income was calculated. Hence, I excluded this household’s budget book. Similarly, a household does not include information on the payments to the credit purchases (household #39). Although this household’s budget book documented the amount purchased in credit purchases, I excluded this household’s budget book because the expenditure series became systematically smaller due to the omission of the payment information.
@@
Regarding the time dimension, I trimmed all the observations before January 12, 1919, and after November 11, 1919, because the number of surveyed households is substantially small in the edge periods. For example, Figure B.4b shows that the number of cross-sectional units is three or less in both periods, causing substantial fluctuations in the net income. Consequently, my analytical sample includes $33$ households measured between January 12 and November 11, 1919. The average number of semi-months per unit is $8.8$ (Std. Dev. $=5.0$). The characteristics of the $33$ THBS households are summarized in detail in Section 3.
@@ B.4 Testing the Potential Influence of the Lack of balance
In this subsection, I test whether attrition leads to selection bias in the THBS sample. If the difference in the preference for consumption predicted large attritions, there must be statistically significant correlations between the family size variables and attrition. In other words, the units with shorter (longer) time-series observations have different preferences from those with longer (shorter) time-series observations. Table B.2 summarizes the results of the balancing tests (Section 3.2). The dependent variable is an indicator variable that takes one if the household has time-series observations less than the median (column 1), 75 percentile (column 2), and 25 percentile (column 3). In Column 1, all the estimated coefficients on the family size variables are close to zero and statistically insignificant, and the Wald statistics $p$-value suggests the null results. This is unchanged if I use the third and first quantiles in columns (2) and (3), respectively. Therefore, the family characteristics are similar between the households with short-term observations and those with long-term observations. This result supports the evidence that the lack of balance in the THBS dataset is unlikely to lead to selection bias.
@@ B.5 Comparing Distributions of the Household Size
Figure B.5 illustrates the household size distributions measured in the census and THBS sample. The Pearson chi-squared test does not reject the null hypothesis of the equality of distributions with $p$-value of $0.948$. This supports the evidence that the THBS sample has a similar household size distribution to that of the entire Tsukishima population (Section 3.1). As described in Section 3.2, while the THBS sample does not cover households with one person and with a very large number of people, it covers most of the classes in the household sizes with reasonable approximation to the population distribution. [^95] The number of households with one people is $362$, accounting for only $6$% of the total number of households. Similarly, the number of households with nine or more people is $294$, accounting for only $5$% (Section 3.2).
@@
Notes: This figure shows the distribution of household size measured in the 1920 Population Census (red) and the THBS sample (green). The Pearson chi-squared test does not reject the null hypothesis of the equality of distributions with $p$-value of $0.948$. The share in each household size measured in the census is used as the theoretical probability in calculating $\chi^{2}$ test statistic (Pearson 1900).
Sources: Created by the author using the Tokyo City Office (1922a, pp. 262–283) and the THBS dataset.
@@ B.6 Estimating Average Head’s Monthly Income
Notes: This figure shows the density of the THBS heads’ age calculated from the THBS daily panel dataset. The dashed line (red) indicates the average age of all the heads (33.2 years old). The dashed line (blue) shows the average age of the heads, excluding an elder head aged 67.6 (32.1 years old). The difference between the two average figures is statistically insignificant with $p=0.4087$.
Source: Created by the author using the THBS dataset.
@@ B.6.1 Daily Wage Distribution by Ages
As the early wage surveys focused only on the average wage for entire workers, official statistics in the 1910s never included average wage statistics by age (bin). In the “Statistical Data Field Survey Law” (Law No. 52) in April 1922, average wages by age group were surveyed for the first time in the “Labor Statistics Field Survey” in October 1924 conducted by the Social Affairs Bureau of the Ministry of Home Affairs (Committee for the History of the Labor Movement 1964, p. 14). The result of this survey in Tokyo was published as the “Labor Statistics Table for Tokyo City and Suburbs” (hereafter, LST) by the Tokyo City Statistics Division in 1926.
@@
Notes: This figure shows the average daily wage distribution by age bins based on the manufacturing census in 1924. Figures B.7a and B.7b show the distributions under 1918 and 1919 prices, respectively. The weighted average daily wage for a 30–39-year-old bin is shown in red. The weighted averages for all and 14+ bins are shown in gray and blue lines, respectively. The fitted fractional polynomial models are shown in the dashed lines. For the 1918 price data, the fitted model is $\hat{y}^{\textit{Daily Wage}}=-10.01+4.15a^{1/2}-0.33a^{2}$ with $R^{2}=0.99$. For the 1919 price data, the fitted model is $\hat{y}^{\textit{Daily Wage}}=-13.18+5.47a^{1/2}-0.43a^{2}$ with $R^{2}=0.99$.
Sources: Created by the author using the Tokyo City Statistics Division (1926a, pp. 16; 244–245). Data on the consumer price index for the cities are obtained from Ohkawa et al. (1967, p. 135).
Based on the LST, Figure B.7 shows the average daily wages of male factory workers in Tokyo City’s machinery and equipment manufacturing industry by age group. In the LST, there are $13$ age bins, and the number of workers and average wages are listed for each bin. Figure B.7a shows the distribution of wages adjusted to 1918 prices using the consumer price index (CPI) for urban areas (Ohkawa et al. 1967, p. 135). The weighted average for ages $14$ and older is $2.58$ yen, which is about the same as the weighted average for daily wages of $2.56$ yen. The $2.58$ yen is between the average values for the 20–24 age bin and the 25–29 age bin. Thus, it is likely to represent the daily wage of workers at approximately $25$. This is consistent with the fact that workers are concentrated in the twenties. However, the weighted average for the 30–39 age bin is $3.02$ yen, which deviates significantly from $2.58$ yen. The distribution of wages at 1919 prices shows a similar result (Figure B.7b). This means that the unweighted average leads to the erroneous conclusion that the sample households are upwardly biased.
@@
where $a$ ($>0$) denotes class value of age bin, $\bm{\varphi}$ is a vector of coefficients and $\mathbf{p}$ is a vector of powers, satisfying $p_{1}<p_{2}<...<p_{m}$. $a^{(p_{j})}$ indicates the Box-Tidwell transformation, taking $\text{ln}(a)$ if $p_{j}=0$ and $a^{p_{j}}$ if $p_{j}\neq 0$. As suggested in Royston and Altman (1994, p. 433), the models requiring a degree higher than two ($m>2$) are rare in practice. The wage by age bin distribution illustrated in Figure B.7 clearly shows a simple quadratic curve. Thus, I consider a fractional polynomial function with degree two as follows:
@@
The maximum likelihood estimation is used to search for the most suitable candidate of the coefficients for a given $m$. Since $m=2$, there are ${}_{8}\mathrm{H}_{2}$ ($=36$) repeated combinations for the candidates of the parameters. By adding the $8$ simple candidates under the nested models (i.e., the models with degree-1 fractional polynomial) to be conservative, I run $44$ regressions in total to find out the model with the highest maximized log-likelihood. The best power vector suggested is $(p_{1},p_{2})=(0.5,1)$. [^97] The deviance, defined as the twice negative (maximized) log-likelihood, is calculated to be $-31.5$. Royston and Altman (1994) proposed the deviance difference, $D(m,\mathbf{p})-D(m,\mathbf{\tilde{p}})$, where $\mathbf{\tilde{p}}$ is the best power vector to compare the different models. This statistic has an asymptotic chi-squared distribution with $m$ degree of freedom. The estimated deviance difference with the simplest linear model (i.e., $m=1$ and $p_{1}=1$) is $54.6$ with $p$-value = $0.000$. Similarly, the difference with an another non-linear model with $m=1$ and $p_{1}=-2$ is $34.2$ with $p$-value = $0.000$. The projection model fitted using the best power vector is as follows:
@@
Next, I shift the predicted distribution based on the difference in the average daily wages in the ASCT and LST. The weighted average of the daily wage is $2.56$ yen for the LST, whereas that for the ASCT is $1.05$ yen, meaning that there is approximately $1.5$ yen difference. [^98] For the LST, I use the number of factory workers in each age bin as weight. The ASCT reports the average daily wage for four different categories of factories in each manufacturing sector: the factory scale (ten+ or below ten workers) and engines (use or not). I calculate the weighted average of the daily wage using the number of factory workers in each age bin (12–14 and 15+) and in each category as weights. Thus, I shift the functional polynomial function by this amount of difference to yield the proxy distribution for 1918.999999This operation, therefore, assumes that the wage-by-age distributions in 1918 and 1924 are reasonably similar. In the LST (i.e., the statistic on which the predicted distribution is based), the weighted average for those aged $14$ and older was $3.25$ yen. In comparison, the overall weighted average is $3.23$ yen, a difference of only $0.02$ yen. A similar trend can be observed in the statistics obtained from the ASCT: the weighted average for those aged $15$ and older is $1.14$, while the overall average is $1.13$, a difference of only $0.01$ yen. This suggests that the wage distribution in the ASCT is similar to that in the LST. Figure B.8a illustrates this operation: dotted curve is a fractional polynomial shifted from the original fractional polynomial shown as a short dashed curve.
@@
The long dashed curve in Figure B.8a examines this relationship by fitting another fractional polynomial model using the 12–14-year old bin figure of the ASCT and the predicted wages for 25–65 from the shifted fractional polynomial.100100100The fitted model is as follows: $\hat{y}^{\textit{Daily Wage}}=-1.19+0.14a^{1}-0.002a^{2},\quad R^{2}=0.98.$ The deviance is reported to be $-149$. I used the 25–65 age range for the prediction based on the age bins used in the LST. Let this curve predict wage distribution for the true distribution. The average daily wage for the 30–40-year-old bin in the shifted fractional polynomial is $1.55$ yen, which is similar to that for the same-age bin in the predicted wage distribution (i.e., $1.52$ yen). However, the wage gaps for the younger ages show more significant differences. If our interest is in the late 20s to the early 40s, it is plausible to employ the predicted wages for the wage comparison in Section 3.
@@
Notes: The short dashed curve illustrates the fractional polynomial (FP) in Figure B.7. The dotted curve indicates the FP shifted by the difference in the average daily wage between the ASCT and LST. The ASCT documents the wage statistics for four different categories for each manufacturing sector: the scale of the factory (10+ or below ten workers) and engines (use or not). I calculated the weighted average of the daily wage using the number of factory workers in each age bin and the category as weights. The long-dashed curve shows the FP refitted using the average daily wage for 12–14-year-old bin ($0.42$ yen) measured in the ASCT and the daily wages for 25–65-year-old suggested in the shifted FP. The average daily wage for a 12–14-year-old bin is only available for the ASCT of 1918. Therefore, this figure for the 1919 price is calculated using the ratio between the average daily wage ($1.05$ yen) and that for 12–14 years old bin ($0.42$) in 1918 as $1.80\times(0.42/1.05)$, where $1.80$ is the average daily wage measured in the ASCT of 1919. The red dashed line is the average daily wage for 30–40-year-old in the shifted FP. The blue dashed line is the average daily wage for 30–40-year-old in the refitted FP.
Sources: Created by the author using the Tokyo City Office (1921, pp. 726–757); Tokyo City Statistics Division (1926a, pp. 16; 244–245).

### v14

@@ B.2 Original Micro-spreadsheets from the Household Budget Survey in Tsukishima Survey
The official report does not provide household-level information. Fortunately, some of the original micro-spreadsheets are preserved in the OISR archives, as Takano was its first director. This study uses the original micro-spreadsheets from the household budget survey (No. 9 in Table B.1), which Gonda used to write the second section of the official report. Although these spreadsheets are not yet publicly available, I obtained permission from the OISR to use the original forms stored in the archives: the OISR, Archives of the Tsukishima Survey (THBS, unreleased). [^76] I would like to express my sincere thanks to Professor Kazue Enoki at the OISR, for allowing me access to the documents. I wish to thank Ms. Mika Nakamura, an archivist at the institute, for her careful assistance during the research.
@@
Notes: This photograph depicts a sample page from the household expenses account of the THBS. The right page documents the details of costs. The left page includes the details of revenue and receiving and making payments on the goods (actual things, including gifts). Source: The THBS, account #6, 15th February, 1919. The tone was adjusted by the author using Adobe Photoshop 24.7.0.
The budget book is titled the Kinsendeiri hikaechō (money receipts and disbursements account book). Its design is based on the form used in Takano’s first household survey (Online Appendix B.1). [^77] Department of Health, Ministry of the Interior (1923a, p. 152). The book is completed daily and comprises four sections: income, expenses, receipts from credit purchases, and receipts and disbursements of actual goods. Most entries appear in the income and expenses sections, indicating that daily life was largely sustained through wages and cash expenses. Figure B.3 provides an example of the budget book.
As noted, the method for selecting households for the survey was not recorded in detail (Online Appendix B.1). According to the report, just over $90$ households were initially recruited through factories, the police station, and elementary schools in Tsukishima. Of these, $50$ households completed the book for at least one month, but $10$ household’s entries were considered incomplete and excluded from the report. In addition to the $40$ budget books used for the report, the OISR holds another set marked as “discarded.” My survey indicates that about six incomplete books from worker households fall into this category. Based on the condition of the remaining books, approximately $50$ households potentially participated in the survey. Among the $40$ households included in the report, two maintained records for more than one year, $11$ households for 6–12 months, and the remaining $27$ households recorded data for less than six months (Department of Health, Ministry of the Interior 1923a, pp. 152–153). Although the survey was initially planned for six months, the average bookkeeping period for these $40$ households was four months.
@@ B.3 Trimming
(c) Monthly panel dataset
@@
Notes: Figure B.4aB.4b, and B.4c show the observations in the daily, adjusted semi-monthly, and adjusted monthly panel datasets, respectively. For example, a marker in Figure B.4a indicates an observation in a specific year-month-date cell. Strictly, semi-monthly, and monthly datasets include both panel and cross-sectional units because aggregation shall trim fractions in the edges of the survey period. Figure B.4b shows 35-panel units and five cross-sectional units (#29; #31-33; #35). Figure B.4c shows 28-panel units and six cross-sectional units (#20; #25; #26; #28; #34; #40). Five cross-sectional units in the semi-monthly panel do not appear in this figure (#29; #31-33; #35). Note that unit #30 having two semi-monthly cells is not included because those cells are separated into different adjusted months (February and March, respectively). Source: Created by the author using the THBS sample.
I digitized all $40$ household budget books. Throughout this digitization process, I found that there are a few misreporting issues in the budget books.
First, there is a household that does not include income information (household #37). Although the official report by Gonda reported the average monthly income for this household (Department of Health, Ministry of the Interior 1923a, p. 183), no documentation enabled me to determine how this average income was calculated. Hence, I excluded this household’s budget book. Similarly, a household does not include information on the payments to the credit purchases (household #39). I excluded this household’s budget book because the expenditure series became systematically smaller due to the omission of the payment information.
@@
Regarding the time dimension, I trimmed all the observations before January 12, 1919, and after November 11, 1919, because the number of surveyed households is substantially small in the edge periods. For example, Figure B.4b shows that the number of cross-sectional units is three or less in both periods, causing substantial fluctuations in the net income. Consequently, my analytical sample includes $33$ households measured between January 12 and November 11, 1919. The average number of semi-months per unit is $8.8$ (Std. Dev. =5.0=5.0). The characteristics of the $33$ THBS households are summarized in detail in Section 3.
@@ B.4 Testing the Potential Influence of the Lack of Balance
If the difference in the preference for consumption predicted large attritions, there must be statistically significant correlations between the family size variables and attrition. In other words, the units with shorter (longer) time-series observations have different preferences from those with longer (shorter) time-series observations. Table B.2 summarizes the results of the balancing tests. The dependent variable is an indicator variable that takes one if the household has time-series observations less than the median (column 1), 75 percentile (column 2), and 25 percentile (column 3). In Column 1, all the estimated coefficients on the family size variables are close to zero and statistically insignificant, and the Wald statistics $p$-value suggests the null results. This is unchanged if I use the third and first quantiles in columns (2) and (3), respectively. Therefore, the family characteristics are similar between the households with short-term observations and those with long-term observations. This result supports the evidence that the lack of balance in the THBS dataset is unlikely to lead to selection bias.
@@ B.5 Comparing Distributions of the Household Size
Figure B.5 illustrates the household size distributions measured in the census and THBS sample. The Pearson chi-squared test does not reject the null hypothesis of the equality of distributions with $p$-value of $0.948$. This supports the evidence that the THBS sample has a similar household size distribution to that of the entire Tsukishima population (Section 3.1). As described in Section 3.2, while the THBS sample does not cover households with one person and with a very large number of people, it covers most of the classes in the household sizes with reasonable approximation to the population distribution. [^78] The number of households with one people is $362$, accounting for only 66% of the total number of households. Similarly, the number of households with nine or more people is $294$, accounting for only 55% (Section 3.2).
@@
Notes: This figure shows the distribution of household size measured in the 1920 Population Census (red) and the THBS sample (green). The Pearson chi-squared test does not reject the null hypothesis of the equality of distributions with $p$-value of $0.948$. The share in each household size measured in the census is used as the theoretical probability in calculating $\chi^{2}$ test statistic (Pearson 1900). Sources: Created by the author using the Tokyo City Office (1922a, pp. 262–283) and the THBS dataset.
@@ B.6 Estimating Average Head’s Monthly Income
Notes: This figure shows the density of the THBS heads’ age calculated from the THBS daily panel dataset. The dashed line (red) indicates the average age of all the heads (33.2 years old). The dashed line (blue) shows the average age of the heads, excluding an elder head aged 67.6 (32.1 years old). The difference between the two average figures is statistically insignificant with $p=0.4087$. Source: Created by the author using the THBS dataset.
@@ B.6.1 Daily Wage Distribution by Ages
As the early wage surveys focused only on the average wage for entire workers, official statistics in the 1910s never included average wage statistics by age (bin). In the Statistical Data Field Survey Law (Law No. 52) in April 1922, average wages by age group were surveyed for the first time in the Labor Statistics Field Survey in October 1924 conducted by the Social Affairs Bureau of the Ministry of Home Affairs (Committee for the History of the Labor Movement 1964, p. 14). The result of this survey in Tokyo was published as the Labor Statistics Table for Tokyo City and Suburbs (hereafter, LST) by the Tokyo City Statistics Division in 1926.
@@
Notes: This figure shows the average daily wage distribution by age bins based on the manufacturing census in 1924. Figures B.7a and B.7b show the distributions under 1918 and 1919 prices, respectively. The weighted average daily wage for a 30–39-year-old bin is shown in red. The weighted averages for all and 14+ bins are shown in gray and blue lines, respectively. The fitted fractional polynomial models are shown in the dashed lines. For the 1918 price data, the fitted model is $\hat{y}^{\textit{Daily Wage}}=-10.01+4.15a^{1/2}-0.33a^{2}$ with $R^{2}=0.99$. For the 1919 price data, the fitted model is $\hat{y}^{\textit{Daily Wage}}=-13.18+5.47a^{1/2}-0.43a^{2}$ with $R^{2}=0.99$. Sources: Created by the author using the Tokyo City Statistics Division (1926a, pp. 16; 244–245). Data on the consumer price index for the cities are obtained from Ohkawa et al. (1967, p. 135).
Based on the LST, Figure B.7 shows the average daily wages of male factory workers in Tokyo City’s machinery and equipment manufacturing industry by age group. In the LST, there are $13$ age bins, and the number of workers and average wages are listed for each bin. The average daily wages have an inverted-U-shape, which is a similar life-cycle pattern to the factory workers observed in late 19th century England (Horrell and Oxley 2000, p. 42). Figure B.7a shows the distribution of wages adjusted to 1918 prices using the consumer price index (CPI) for urban areas (Ohkawa et al. 1967, p. 135). The weighted average for ages $14$ and older is $2.58$ yen, which is about the same as the weighted average for daily wages of $2.56$ yen. The $2.58$ yen is between the average values for the 20–24 age bin and the 25–29 age bin. Thus, it is likely to represent the daily wage of workers at approximately $25$. This is consistent with the fact that workers are concentrated in the twenties. However, the weighted average for the 30–39 age bin is $3.02$ yen, which deviates significantly from $2.58$ yen. The distribution of wages at 1919 prices shows a similar result (Figure B.7b). This means that the unweighted average leads to the erroneous conclusion that the sample households are upwardly biased.
@@
where $a$ (>0>0) denotes class value of age bin, $\bm{\varphi}$ is a vector of coefficients and $\mathbf{p}$ is a vector of powers, satisfying $p_{1}<p_{2}<...<p_{m}$. $a^{(p_{j})}$ indicates the Box-Tidwell transformation, taking $\text{ln}(a)$ if $p_{j}=0$ and $a^{p_{j}}$ if $p_{j}\neq 0$. As suggested in Royston and Altman (1994, p. 433), the models requiring a degree higher than two ($m>2$) are rare in practice. The wage by age bin distribution illustrated in Figure B.7 clearly shows a simple quadratic curve. Thus, I consider a fractional polynomial function with degree two as follows:
@@
[…]
@@
Next, I shift the predicted distribution based on the difference in the average daily wages in the ASCT and LST. The weighted average of the daily wage is $2.56$ yen for the LST, whereas that for the ASCT is $1.05$ yen, meaning that there is approximately $1.5$ yen difference. [^81] For the LST, I use the number of factory workers in each age bin as weight. The ASCT reports the average daily wage for four different categories of factories in each manufacturing sector: the factory scale (ten+ or below ten workers) and engines (use or not). I calculate the weighted average of the daily wage using the number of factory workers in each age bin (12–14 and 15+) and in each category as weights. Thus, I shift the functional polynomial function by this amount of difference to yield the proxy distribution for 1918.828282This operation, therefore, assumes that the wage-by-age distributions in 1918 and 1924 are reasonably similar. In the LST (i.e., the statistic on which the predicted distribution is based), the weighted average for those aged $14$ and older was $3.25$ yen. In comparison, the overall weighted average is $3.23$ yen, a difference of only $0.02$ yen. A similar trend can be observed in the statistics obtained from the ASCT: the weighted average for those aged $15$ and older is $1.14$, while the overall average is $1.13$, a difference of only $0.01$ yen. This suggests that the wage distribution in the ASCT is similar to that in the LST. Figure B.8a illustrates this operation: dotted curve is a fractional polynomial shifted from the original fractional polynomial shown as a short dashed curve.
@@
The long dashed curve in Figure B.8a examines this relationship by fitting another fractional polynomial model using the 12–14-year old bin figure of the ASCT and the predicted wages for 25–65 from the shifted fractional polynomial. [^83] The fitted model is as follows: $\hat{y}^{\textit{Daily Wage}}=-1.19+0.14a^{1}-0.002a^{2},\quad R^{2}=0.98.$ The deviance is reported to be 149-149. I used the 25–65 age range for the prediction based on the age bins used in the LST. Let this curve predict wage distribution for the true distribution. The average daily wage for the 30–40-year-old bin in the shifted fractional polynomial is $1.55$ yen, which is similar to that for the same-age bin in the predicted wage distribution (i.e., $1.52$ yen). However, the wage gaps for the younger ages show more significant differences. If our interest is in the late 20s to the early 40s, it is plausible to employ the predicted wages for the wage comparison in Section 3.
@@
Notes: The short dashed curve illustrates the fractional polynomial (FP) in Figure B.7. The dotted curve indicates the FP shifted by the difference in the average daily wage between the ASCT and LST. The ASCT documents the wage statistics for four different categories for each manufacturing sector: the scale of the factory (10+ or below ten workers) and engines (use or not). I calculated the weighted average of the daily wage using the number of factory workers in each age bin and the category as weights. The long-dashed curve shows the FP refitted using the average daily wage for 12–14-year-old bin ($0.42$ yen) measured in the ASCT and the daily wages for 25–65-year-old suggested in the shifted FP. The average daily wage for a 12–14-year-old bin is only available for the ASCT of 1918. Therefore, this figure for the 1919 price is calculated using the ratio between the average daily wage ($1.05$ yen) and that for 12–14 years old bin ($0.42$) in 1918 as $1.80\times(0.42/1.05)$, where $1.80$ is the average daily wage measured in the ASCT of 1919. The red dashed line is the average daily wage for 30–40-year-old in the shifted FP. The blue dashed line is the average daily wage for 30–40-year-old in the refitted FP. Sources: Created by the author using the Tokyo City Office (1921, pp. 726–757); Tokyo City Statistics Division (1926a, pp. 16; 244–245).